OpenStudy (anonymous):

Change to standard form.y^2-4x^2-16x-2y-19=0

OpenStudy (anonymous):

y^2-4x^2-16x-2y-19=0 so put the x together and the y together and the constant the last in the equation. so it will be y^2-2y-4x^2-16x-19=0 then factorise; y(y-2)-4x(x+4)-19=0